At present A is five times older than B. After 12 years A would be twice as old as B then what are their ages at present
Answer by Cromlix(4381)

Hi there,
A is five times older than B
A = 5B
After 12 years A would be twice as old as B
A + 12 = 2(B + 12)
Replace A with 5B from first equation
5B + 12 = 2(B + 12)
Multiply out
5B + 12 = 2B + 24
Collect like terms
5B - 2B = 24 - 12
3B = 12
B = 4
A is 20 yrs old
B is 4 yrs old
Hope this helps:-)